CTA deployment meeting

600/R-001 (CERN)



Show room on map
Michael Davis (CERN)
    • 14:00 14:10
      Getting LHCb into Production 10m


      • Transfers from EOS LHCb to EOSCTA LHCb are completing successfully following EOS/XRootD upgrade.
      • DAQ tests (including file safely on tape check) will be tested after the migration, in production.
      • All that remains to test before migration is grid transfers to/from T1s.


      • w/c 22 February: final tests. Prepare OTGs.
      • 1 March: disable write access to CASTOR LHCb in preparation for migration. Check all queues are flushed and everything is written to tape.
      • 8 March: disable CASTOR LHCb. Migrate to CTA.
      • 15 March: EOSCTA LHCb in production

      To Do

      • HTTP+TPC tests
      • Publish OTGs
    • 14:10 14:20
      Getting PUBLIC into Production 10m

      Status Update

      • Spinner space has been configured (#161)
      • AMS tests: problem fixed, waiting for Baosong to resume tests
      • NA61/SHINE: Met last week with EOS+FTS teams to discuss DAQ storage/best practices. Eddie is following up.
      • DUNE. No news, but we are not pushing. See #213
      • n_TOF will repeat their CASTOR stress test with CTA
      • TOTEM meeting on Thursday 15:00
      • Repacking LEP-era data to improve tape occupancy in preparation for migration

      TO DO

      • Test garbage collector on spinner space
      • n_TOF, COMPASS: Recall tests with spinner space
      • Finish repack of public_user
      • Publish Knowledge Base article for users to access files in CTA, see #214 and draft article in CodiMD
      • Migrate data from legacy experiments (Aleph, Chorus, Delphi, Nomad, Opal, ...)
      • Bartek (NA61) is following up on issue of experiment data in CASTOR /user part of the namespace


      • w/c 22 February: n_TOF DAQ stress test on EOSCTA PUBLIC PPS, NA62 stress test on EOSCTA PPS (readonly)
      • 8 March: IT/SME coordination meeting
      • w/c 15 March: migrate NA62 to CTA
    • 14:20 14:30
      CTA Software 10m
      • #959 Cedric's departure preparation
      • Review of any work we would like Cedric to finish off
    • 14:30 14:35
      AOB 5m